# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v9

@@ 3.1 Bermudan Option with $d=1$ under Black-Scholes
[…]
@@
However, the traditional LSM is showing poor agreement with discrepancy up to $10\%+$ for high spot; for low spots below 55% and high spots above 135%, the fitted curve is even below the exercise payoff $Z(S)$ and thus arbitragable, which mistakenly instructs the option holder to exercise immediately. Apart from mis-fitting away from ATM, another cause of poor results is that at $t_{k=30}=\frac{T}{2}$ the numerical errors have been accumulated after 29 LSM iterations from maturity. Nevertheless, FD-LSM doesn’t suffer from these two issues thanks to the use of the FD solution as the ansatz to be regressed on.
@@ 3.4 EPE/CVA for European option in $d=4$ under Black-Scholes
Since the portfolio has only one positive payoff to the party, it follows that its discounted option value under $\operatorname*{\mathbb{Q}}$ measure is a martingale, hence its expected value, i.e. EPE*, is constant over time:
@@ 3.5 WIC from $d=5$ to $d=50$ under Black-Scholes and Local Volatility
# $d$ $\rho$ $T$ $\frac{1}{\Delta T}$ $c$ $r$ LSM ($R$=4) FD-LSM ($R$=4) PV $\frac{\text{PV Diff}}{\text{s.e.}}$ c.t. Diff%
@@
# $d$ $\rho$ $T$ $\frac{1}{\Delta T}$ $c$ $r$ LSM ($R$=4) FD-LSM ($R$=4) PV $\frac{\text{PV Diff}}{\text{s.e.}}$ c.t. Diff%
@@ A.4 Local Regression Error in LSM
$\begin{split}&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}]\\ =&\operatorname{\mathbb{E}}[(\Delta\vec{\beta}^{P^{\text{LSM}}}_{k}\cdot\vec{\phi}^{P^{\text{LSM}}}_{R,k})^{2}]\\ =&\operatorname{\mathbb{E}}[\text{tr}(\Delta\vec{\beta}^{P^{\text{LSM}}T}_{k}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\Delta\vec{\beta}^{P^{\text{LSM}}T}_{k}\vec{\phi}^{P^{\text{LSM}}}_{R,k})]\\ =&\operatorname{\mathbb{E}}[\text{tr}(\vec{\phi}^{P^{\text{LSM}}T}_{R,k}(e^{-r\Delta T}V_{k+1}-F_{k})^{2}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k})]\qquad\qquad\qquad\qquad\quad\text{ (Law of Iterated Expectation)}\\ =&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}\text{tr}(\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1})]\qquad\qquad\qquad\qquad\qquad\qquad\text{(Cyclic property of matrix trace)}\\ \leq&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}]\cdot\operatorname{\mathbb{E}}[\text{tr}(\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1})]\qquad\qquad\qquad\qquad\qquad\qquad\text{(Cauchy-Schwarz inequality)}\\ \leq&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}]\cdot\text{tr}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}](\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}]\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1})\qquad\qquad\qquad\qquad\qquad\qquad\text{ (Cauchy-Schwarz inequality)}\\ =&\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})]\cdot\text{tr}({\bf{I}}_{R})\\ =&\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})]\cdot R,\end{split}$ (4.3)
@@ A.6.1 Inductive Proof of Convergence under finite $M$
A.6.2 Fixed Point Analysis in the limit of $M\to\infty$
The asymptotic behavior of $x_{k}$ in the limit of $k\to 1$ and $M\to\infty$ depends on $a_{0}=\lim\limits_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}a_{k}$ and $c_{0}=\lim\limits_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}c_{k}$.

### v10

@@ 3.1 Bermudan Option with $d=1$ under Black-Scholes
We will compare the option $PV$ under different approaches: PDE1D, LSM, $\text{LSM}^{\text{CV}}$, FD-LSM and $\text{FD-LSM}^{\text{CV}}$ as discussed in Sec.2. The superscript (CV) indicates the use of 1d analytic European option values as the control variates in pricing stage (Tian and Burrage, (2003)).
@@
However, the traditional LSM is showing poor agreement with discrepancy up to $10\%+$ for high spot; for low spots below 55% and high spots above 135%, the fitted curve is even below the exercise payoff $Z(S)$ and thus arbitragable, which mistakenly instructs the option holder to exercise immediately. Apart from mis-fitting away from ATM, another cause of poor results is that at $t_{k=30}=\frac{T}{2}$ the numerical errors have been accumulated after 30 LSM iterations from maturity. Nevertheless, FD-LSM doesn’t suffer from these two issues thanks to the use of the FD solution as the ansatz to be regressed on.
@@ 3.4 EPE/CVA for European option in $d=4$ under Black-Scholes
[…]
@@ 3.5 WIC from $d=5$ to $d=50$ under Black-Scholes and Local Volatility
# $d$ $\rho$ $T$ $\frac{1}{\Delta T}$ $c$ $r$ LSM (RR=4) FD-LSM (RR=4) $\frac{\text{PV Diff}}{\text{s.e.}}$ c.t. Diff%
@@
# $d$ $\rho$ $T$ $\frac{1}{\Delta T}$ $c$ $r$ LSM (RR=4) FD-LSM (RR=4) $\frac{\text{PV Diff}}{\text{s.e.}}$ c.t. Diff%
@@ A.4 Local Regression Error in LSM
[…]
@@ A.6.1 Inductive Proof of Convergence under finite $M$
A.6.2 Fixed Point Analysis in the limit as $M\to\infty$
The asymptotic behavior of $x_{k}$ in the limit as $k\to 1$ and $M\to\infty$ depends on $a_{0}=\lim\limits_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}a_{k}$ and $c_{0}=\lim\limits_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}c_{k}$.